OpenStudy (anonymous):

find f of each given value of x: f(x)=3x^2+4x-2

OpenStudy (anonymous):

wat r the given values of x ??

OpenStudy (anonymous):

replace x by values

OpenStudy (hba):

@princess_a7 Your question is incomplete mam.

OpenStudy (anonymous):

that's what it says.. find f of each given value of x.

OpenStudy (hba):

@princess_a7 Does It Have Any Thing Like f(1) f(-6) ?

OpenStudy (hba):

As in Parts ?

OpenStudy (hba):

@princess_a7 Because the information is not sufficient for me to answer your question :)

OpenStudy (anonymous):

nope it just gives me the problem: f(x)=3x^2+4x-2

OpenStudy (anonymous):

\[3x ^{2}+4x-2 = (x+ \frac{2}{3})^{2}-\frac{10}{9}\]

OpenStudy (hba):

Ok you just have to plug in whatever value you get in place of x,for example lets suppose we had to find f(-3) So, f(x)=3x^2+4x-2 f(x)=3(-3)^2+4(-3)-2 f(x)=27-12-2 f(x)=13

OpenStudy (hba):

@amir.sat This is functions.

OpenStudy (anonymous):